Dr. David Bryan Carpenter

Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244-4100 Phone: (315) 443-5068 FAX : (315) 443-1973 email: dbc@npac.syr.edu

February 29, 2000

Departure of Sang Boem Lim

Dear Housing Office,

Sang Boem Lim is a Ph.D. student working under my supervision in the University’s Northeast Parallel Architecture Center. The Center will be closing imminently due to the departure of its director, Professor Geoffrey Fox, for Florida State University. To continue his studies, Sang Lim has agreed to follow Professor Fox and me to Florida.

I hope you can release him from the lease agreement covering his tenancy in Slocum Heights, the circumstances being essentially outside his control. Currently we expect to move at the end of March.

Yours sincerely,

Bryan Carpenter